If someone could help on this it would be greatly appreciated.
For the systems of linear equation
-Determine how many solutions exist
-Use either elimination or substitution to find the solutions
-Graph the two lines labeling the x intercepts, y intercepts, and points of intersection.
x + y = 3 and y = x + 3

Answer by ewatrrr(10682)

 
Hi,
x + y = 3 and
y = x + 3 OR -x+y = 3
x + y = 3 GREEN (3,0) x-intercept (0,3) y-intercept
-x + y = 3 BLUE (-3,0) x-intercept (0,3) y-intercept
2y = 6
y = 3 and x = 0 |||Note: (x+y=3)
Pt(0,3) is the Pt of Intersection